Vapour density of moxture of NO2 and N2O4 is 34.5, then % abundance of NO2 in mixture is ______?


a. 50%


b. 25%


c. 40%


d. 60%

Grade:11

1 Answers

Yogita Nain
12 years ago

Let total moles of mixture is 100.

 

let there be x moles of NO2 molecular mass is 46 g

 

thus there are (100-x) moles of N2O4 molecular mass is 92g .  

 

total mass of mixture =  x * 42 + (100-x) 92  

 

molecular mass of mixture  =  2 * V.D = 2* 34.5 = 69

(x*42 + (100-x)92) / 100 = 69

 

solving... x=46

percentage abundance = ( (46*42) / (46*42 + 64*92) ) *100 = 24.7 (approx 25)